QUESTION 4 10 points Save Answer A company buys $62,062 of computers. What is the depreciation for this asset in year 3?
Answer #1

Solution:-

Computer is depreciated under property class of Five year property in Modified Accelerated Cost Recovery System(MACRS).

Years Depreciation(%) Depreciation Charge
1 20.00 $62,062 * 20 / 100 = 12,412.4
2 32.00 $62,062 * 32 / 100 = 19,859.84
3 19.20 $62,062 * 19.20 / 100 = 11,915.90
4 11.52 $62,062 * 11.52 / 100 = 7,149.54
5 11.25 $62,062 * 11.52 / 100 = 7,149.54
6 5.76 $62,062 * 5.76 / 100 = 3,574.77

According to MACRS the depreciation for year 3 is $ 11,915.90.